OpenStudy (anonymous):

Will give medal!!!! Find the slope and y-intercept- 7. y=3x+1 9. y=2x-5 11. y=5x-3 13. y=4 15. y=1/4x-1/3 @magepker728

OpenStudy (anonymous):

@wizfan @WHAT?! @Abhisar

OpenStudy (anonymous):

y=mx+b m is slope b is y-intercept

OpenStudy (anonymous):

y=3x+1 slope=3 y-intercept=1

OpenStudy (anonymous):

7.m=3/1 and y=1

OpenStudy (anonymous):

slop is not 3 but 3/1

OpenStudy (anonymous):